A bag contains 2 A's, 3 B's, and 4 C's. You choose a letter and replace it. Then you choose another. What is the probability of choosing the combinations below:
1. (A)
2. (A,B)
3. (B,B)
4. (C,C)
5. (B,C)
6. (C,B)

Answer by sudhanshu_kmr(1152)

1. (A) probability = 2/9

2. (A,B) probability = 2/9 *3/9 = 2/27

3. (B,B) probability = 3/9 *3/9 = 1/9

4. (C,C) probability = 4/9 * 4/9 = 16/81

5. (B,C) probability = 3/9 * 4/9 = 4/27

6. (C,B) probability = 4/9 * 3/9 = 4/27